# Definability of band structures on posets

### v1

@@ 5 Conclusion
We have presented many examples which showcase the usefulness of an order theoretical point of view for studying bands. We proved that we can define, for certain posets, a band operation by invoking some of its structural properties, such as the existence of a special order-preserving function for a normal poset, or a decomposition into disjoint convex subchains with minimum for a foliated tree. More so, we showed that in the first case, this is the only way of defining a right-normal band operation over a normal poset. Here, natural definability questions arise. There appears to be no way of canonically assigning a right-normal band operation to every normal poset. We can then ask ourselves if being able to define this assignment is equivalent to the Axiom of Choice or some fragment of it.
We proved the equivalence of the associativity of foliated trees and the Axiom of Choice. This, together with the observations regarding right-normal bands, shows that assigning a (right-regular or right-normal) band operation to an associative poset is highly non-canonical. We believe that a broader family of trees might be proven to be associative. It appears to be that for associative trees there must exist, perhaps as some sort of limit, some order type that “occurs densely”. That is, a type order which appears in arbitrary low levels of the tree. Therefore, this leads us to believe that every tree admiting a decomposition into disjoint convex subchains with isomorphic initial segments might be associative.
The following is a list of problems/questions which remain open:
Question 5.1.
@@
Question 5.2.
@@
A result of this sort would be an analogue of Theorem 2.25
Question 5.3.
@@
We believe the answer to this question to be negative. From Theorem 2.29 we know that every disjoint union of associative posets with minimum is associative, so if a counterexample exists, at leats one of the associative posets must not have a minimum.
Question 5.4.
@@
Question 5.5.
Is there a more suitable notion of morphism of associative posets which yields a more interesting category?
@@ Acknowledgment
: We thank Dr. Miguel Campercholi for suggesting that we study the existence of a left adjoint for the forgetful functor between the categories of RRB and associative poset.
@@ Appendix A Additional proofs
A.1 The crown is not associative
@@
Assume it is related to a posemigroup. Then all previous lemmas apply.
@@
A.2 Preimages of foliated trees
@@
As $(x\mathbin{\tilde{\cdot}}y)<z$ by Claim  5.
@@
because $(x\mathbin{\tilde{\cdot}}y)$ must be incomparable with $z$ by Claim 6 as $x\mathbin{\tilde{\cdot}}y\leq y$.

### v2

@@ 7 Conclusion
We have presented many examples which showcase the usefulness of an order theoretical point of view for studying bands. We proved that we can define, for certain posets, a band operation by invoking some of its structural properties, such as the existence of a special order-preserving function for a normal poset, or a decomposition into disjoint convex sub-chains with minimum for a foliated tree. More so, we showed that in the first case, this is the only way of defining a right-normal band operation over a normal poset. Here, natural definability questions arise. There appears to be no way of canonically assigning a right-normal band operation to every normal poset. We can then ask ourselves if being able to define this assignment is equivalent to the Axiom of Choice or some fragment of it.
We proved the equivalence of the associativity of foliated trees and the Axiom of Choice. This, together with the observations regarding right-normal bands, shows that assigning a (right-regular or right-normal) band operation to an associative poset is highly non-canonical. We believe that a broader family of trees might be proven to be associative. It appears to be that for associative trees there must exist, perhaps as some sort of limit, some order type that “occurs densely”. That is, a type order which appears in arbitrary low levels of the tree. Therefore, this leads us to believe that every tree admitting a decomposition into disjoint convex sub-chains with isomorphic initial segments might be associative.
The following is a list of problems/questions which remain open.
Question 7.1.
@@
Regarding the definability of RNB structures over normal posets, we have:
Question 7.2.
@@
A result of this sort would be an analogue of Theorem 6.2. We are also interested in a weaker version of the previous question:
Question 7.3.
The following dwells on possible extensions of Lemma 4.2 or Lemma 4.3:
Question 7.4.
@@
We believe the answer to this question to be negative. From Theorem 6.8 we know that every disjoint union of associative posets with minimum is associative, so if a counterexample exists, at least one of the associative posets must not have a minimum.
Question 7.5.
@@ Acknowledgment
We warmly thank the referee for their valuable feedback that greatly helped to improve the paper, especially the pointer to Green’s work and suggestions to our naming conventions.
@@ Appendix A Additional proofs
A.1 Multiplication tables
We present here the multiplication tables of some of the examples presented. In particular, those shown in Section $3$. We begin giving the multiplication table for the RNB mentioned in Example 3.5.
Consider the following labeling on the poset:
[…]
The multiplication table of the RNB obtained is:
$\cdot$ $a$ $b$ $c$ $d$ $e$
$a$ $a$ $b$ $c$ $d$ $e$
$b$ $a$ $b$ $c$ $d$ $e$
$c$ $c$ $c$ $c$ $e$ $e$
$d$ $d$ $d$ $e$ $d$ $e$
$e$ $e$ $e$ $e$ $e$ $e$
We will now present the multiplication tables for each of the operations obtained in Example 3.6. Consider the following labeling on the poset:
[…]
The multiplication tables of the RNBs corresponding to the previous example are, respectively:
$\cdot$ $a$ $b$ $c$ $d$ $e$
$a$ $a$ $e$ $c$ $e$ $e$
$b$ $e$ $b$ $e$ $d$ $e$
$c$ $c$ $e$ $c$ $e$ $e$
$d$ $e$ $d$ $e$ $d$ $e$
$e$ $e$ $e$ $e$ $e$ $e$
$\cdot$ $a$ $b$ $c$ $d$ $e$
$a$ $a$ $d$ $c$ $d$ $e$
$b$ $c$ $b$ $c$ $d$ $e$
$c$ $c$ $d$ $c$ $d$ $e$
$d$ $c$ $d$ $c$ $d$ $e$
$e$ $e$ $e$ $e$ $e$ $e$
$\cdot$ $a$ $b$ $c$ $d$ $e$
$a$ $a$ $b$ $c$ $d$ $e$
$b$ $a$ $b$ $c$ $d$ $e$
$c$ $c$ $d$ $c$ $d$ $e$
$d$ $c$ $d$ $c$ $d$ $e$
$e$ $e$ $e$ $e$ $e$ $e$
A.2 Every poset having at most $4$ elements is associative
Here we will show that every poset having at most $4$ elements is associative.
  • Posets having at most $3$ elements are either a meet-semilattice or a forest of foliated trees (Remark 6.7).
  • All but two posets with $4$ elements are either a meet-semilattice, a forest of foliated trees or an ordered sum of antichains (Lemma 4.1).
This leaves us to prove the associativity of only two $4$-element posets. These are:
$\bullet$00$\bullet$1111$\bullet$2222$\bullet$3333 \$b$
For the first poset, note that we can prove its associativity using Theorem 6.8. Using this, we can provide a RRB structure given by:
$\cdot$ 00 $1$ $2$ $3$
00 00 $3$ $3$ $3$
$1$ 00 $1$ $3$ $3$
$2$ 00 $3$ $2$ $3$
$3$ 00 $3$ $3$ $3$
As for the second poset, the only admissible RRB is given by:
$\cdot$ $a$ $b$ $c$ $d$
$a$ $a$ $c$ $c$ $d$
$b$ $c$ $b$ $c$ $d$
$c$ $c$ $c$ $c$ $d$
$d$ $c$ $d$ $c$ $d$
A.3 The crown is not associative
@@
Assume it is related to a right posemigroup. Then all previous lemmas apply.
@@
A.4 Preimages of foliated trees
@@
As $(x\mathbin{\tilde{\cdot}}y)<z$ by Claim  6.9.
@@
because $(x\mathbin{\tilde{\cdot}}y)$ must be incomparable with $z$ by Claim 6.10 as $x\mathbin{\tilde{\cdot}}y\leq y$.